A silver ring is composed of 1.02×1023 atoms.1.02×1023 atoms. Calculate the mass of the ring in grams.
Answer #1

1 mol of silver contains mass of 107.8 g.

We know ,

6.022 \times 1023 atoms of silver = 1 mol of silver

=> 1 atoms of silver = (1\div 6.022 \times 1023) mol of silver

=> 1.02 \times 1023  atoms of silver = (1.02 \times 1023\div 6.022 \times 1023) mol of silver = (1.02\div6.022) mol of silver

1 mol of silver contains mass of 107.8 g.

=>(1.02\div6.022) mol of silver contains mass of = 107.8 \times (1.02\div6.022) g = 18.26 g